# neurips_2025.bib

### v1

@@ 4.1 Toy Examples
FNO 0.81$\pm$0.17 69.5$\pm$8.2
Unet 0.52$\pm$0.35 125.7$\pm$8.5
FNO($\frac{d\mathbf{u}}{dt}$) 0.044$\pm$0.002 75.5$\pm$1.9
Unet($\frac{d\mathbf{u}}{dt}$) 0.068$\pm$0.029 127.4$\pm$4.9
HNF 0.0039$\pm$0.0002 150.9$\pm$3.3
Table 2: Results for 1D PDEs. Parameter counts are: FNO (65K), Unet (65K), HNF (32K) for Adv, and FNO (135K), Unet (146K), HNF (87K) for KdV.
Figure 3: Hamiltonians for Adv and KdV for predicted trajectories, plotted over time. FNO and Unet models are shown with their $d\mathbf{u}/dt$ variants. HNFs can better conserve Hamiltonians and remain stable over time.
Two baselines are considered, a FNO [li2021fourierneuraloperatorparametric] and a Unet [gupta2022multispatiotemporalscalegeneralizedpdemodeling] model. Additionally, the HNF model is used with a nonlinear, FiLM-conditioned SIREN kernel. Global conditioning is used in the KdV experiments by calculating the scale/shift parameters as $\gamma_{i}^{(l)}=[\gamma_{\theta}^{(l)}(\mathbf{u}^{t})]_{i}$ and $\beta_{i}^{(l)}=[\beta_{\theta}^{(l)}(\mathbf{u}^{t})]_{i}$; the global FiLM networks are each parameterized by a shallow, 1D CNN. We find that global conditioning is needed in more complex PDEs since the kernel output $z_{i}=k_{\theta}(\cdot)$ can depend on more than just $u_{i}$ at a given point $x_{i}$. In particular, the Hamiltonian for the KdV equation has a $u_{xx}$ term that requires non-local knowledge of $\mathbf{u}^{t}$. Lastly, to ensure a fair comparison, we also train variants of the FNO and Unet baselines that predict temporal derivatives $\text{d}\mathbf{u}/\text{d}t$ [Zhou_2025], as opposed to the usual variants that predict $\mathbf{u}^{t+1}$; these use the same numerical integrator during inference as the HNF model.
The metrics that are reported are the rollout error = $\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}(\mathbf{u}^{t},\mathbf{u}_{\theta}^{t})$, and the correlation time = $\max(t)\>s.t.\>\mathcal{C}(\mathbf{u}^{t},\mathbf{u}_{\theta}^{t})>0.8$. The loss used is a relative L2 loss, or $\mathcal{L}(\mathbf{u}^{t},\mathbf{u}_{\theta}^{t})=\frac{||\mathbf{u}^{t}-\mathbf{u}^{t}_{\theta}||_{2}^{2}}{||\mathbf{u}^{t}||_{2}^{2}}$ and the correlation criterion $\mathcal{C}$ is the Pearson correlation. Rollout error is used to measure the accuracy of the predicted trajectory; for chaotic dynamics rollout error is often skewed after autoregressive drift, therefore correlation time measures the portion of the trajectory that is accurate. Each metric is averaged over the validation set and the mean and standard deviation over three seeds is reported in Table 2. Furthermore, given a predicted trajectory $[\mathbf{u}_{\theta}^{1},\ldots,\mathbf{u}_{\theta}^{T}]$ from a model, we can examine its Hamiltonian at each timestep and plot the scalar value $\mathcal{H}[\mathbf{u}_{\theta}^{t}]$ over time, shown in Figure 3.
We observe that the proposed HNF model is able to outperform current baselines, both on quantitative metrics and when conserving Hamiltonians over time. HNFs are also efficient, capable of achieving better performance with nearly half the parameters of baseline models due to sharing weights between input points $x_{i},u_{i}$. Furthermore, despite training on a shorter time horizon, HNFs can extrapolate and remain stable in longer prediction horizons. This is more pronounced in the KdV equation, where initial states are smoother than later, more chaotic states. An example trajectory is shown in Figure 5, where the baselines struggle to predict later states since these are not seen during training.
@@
To evaluate HNFs on a more complex system, we consider the 2D Shallow Water Equations (SWE). These equations are an approximation of the Navier-Stokes equations when the horizontal length scale is much larger than the vertical length scale. We consider its conservative form, which is often used as a simple model for atmosphere or ocean dynamics.
For vector-valued functions $\mathbf{u}(x,y)\in\mathbb{R}^{d}$, the operator $\mathcal{J}$ becomes a $d\times d$ matrix of operators. For the 2D Shallow Water Equations, the operator matrix $\mathcal{J}$ is given by [ALLEN1996229, Ripa1993, Dellar2003]:
@@
Data for the 2D SWE system is generated from PyClaw [clawpack, pyclaw-sisc, mandli2016clawpack] using random sinusoidal initial conditions (Sines) for the height:
@@
FNO 0.057$\pm$0.002 0.117$\pm$0.0009
Unet 0.010$\pm$0.0014 0.042$\pm$0.0006
HNF 0.026$\pm$0.0003 0.021$\pm$0.0015
Table 3: Rollout errors for 2D SWE. Parameter counts are: FNO (7M), Unet (3M), HNF (3M). After training, models are evaluated on in-distribution ICs (Sines) or out-of-distribution ICs (Pulse).
Figure 4: Hamiltonians for Sines and Pulse ICs, plotted over time. Despite potentially larger rollout errors, HNFs can better conserve the Hamiltonian of predicted solutions.
@@
We again consider benchmarking against Unet and FNO models for 2D systems, and use a global, FiLM-conditioned SIREN kernel for the HNF model. Validation rollout errors for in-distribution ICs (Sines) and out-of-distribution ICs (Pulse) are given in Table 3, where the mean and standard deviation is calculated over three seeds. Additionally, the Hamiltonian over time for model rollouts on these test sets are plotted in Figure 4. Interestingly, the derivative variants of the baseline models could not train stably, possibly due to sharp gradients that develop in the absence of viscosity. Additionally, the baselines are trained with the benefit of normalization, as the height channel has a different scale than the velocity channels; however, since the Hamiltonian is not invariant under scaling and shifting, normalization was not performed for the HNF model.
@@
For test samples within the training distribution (Sines), conventional neural surrogates are able to perform well and achieve low loss, and HNFs are on-par with these models. However, when examining the Hamiltonian, HNFs can predict solutions that better conserve physical invariances. We validate this in Appendix B.1 by calculating the error of the Hamiltonian of predicted trajectories across different experiments. This ability to conserve the Hamiltonian helps HNFs to generalize to unseen initial conditions, and it achieves better performance on the Gaussian Pulse test set. Interestingly, it also conserves the Hamiltonian on out-of-distribution inputs, suggesting a strong inductive bias in the model. This is an empirically observed bias, to understand what mechanisms enable energy conservation, we ablate different methods within HNFs to understand their contributions, shown in Appendix B.1. Lastly, to visualize model performance, predicted trajectories of Sines and Pulse ICs are shown in Figures 6 and 7.
5 Discussion
@@
Despite an interesting theoretical basis and its novelty, HNFs have some limitations that we hope future work can address. Firstly, the theory for nonlinear functionals is underdeveloped. Another limitation is the additional overhead during inference to backpropagate through the network and evaluate the linear operator $\mathcal{J}$. We quantify this in section B.2 in the Appendix. Additionally, implementing $\mathcal{J}$ without a neural operator may require care to avoid numerical errors; this is further discussed in Appendix D.1. Furthermore, HNFs are only applicable to Hamiltonian systems; this would exclude systems with viscosity or friction, although modifications exist to extend Hamiltonian mechanics to non-conservative systems [sosanya2022dissipativehamiltonianneuralnetworks, hh_decomp]. Even for conservative PDE systems, implementing a Hamiltonian structure is not as straightforward as training a neural surrogate with next-step prediction. Lastly, while not a limitation, an interesting observation is that introducing a loss on $\mathcal{H}[\mathbf{u}^{t}]-\mathcal{H}[\mathbf{u}_{\theta}^{t}]$ does not improve HNF performance; we hypothesize that since the Hamiltonian is conserved over time, learning the identity mapping is the easiest way to minimize this error and thus degrades model training.
@@
Despite these limitations, neural functionals represent a new class of models for approximating functionals and their derivatives. Within a Hamiltonian framework, HNFs show good performance in predicting PDE solutions and have excellent capabilities in conserving energy-like quantities. Interestingly, molecular dynamics (MD) is a highly promising use case; on the atomistic scale, energies are always conserved and its derivative (forces) are used to evolve simulations. For ML-based MD simulations, conserving energy is highly important due to extremely long rollouts, as timesteps $\Delta t$ are on the picosecond scale [fu2023forcesenoughbenchmarkcritical]. Apart from MD, functionals arise naturally in many physical systems. While the applications are still speculative, functionals are a concise and meaningful description of a physical state, such as describing the drag or weight of a continuum system, and having meaningful functional derivatives may have interesting uses for optimization.
6 Related Works
To the best of the authors’ knowledge, this is the first time an architecture has been developed to learn function to scalar mappings. Therefore, we focus on discussing related works in PDE surrogate modeling and neural networks with Hamiltonian structure. Original work on Hamiltonian Neural Networks (HNNs) [greydanus2019hamiltonianneuralnetworks, onlearninghamiltonian] established them as a promising architecture to respect physical laws, while later work improved their performance [DAVID2023112495, HORN2025113536, Zhong2020Symplectic, Chen2020Symplectic, sanchezgonzalez2019hamiltoniangraphnetworksode], investigated the mechanisms behind HNNs [gruver2022deconstructinginductivebiaseshamiltonian], or used HNNs in different applications [Mattheakis_2022, Toth2020Hamiltonian, zhong2021extendinglagrangianhamiltonianneural]. A closely related set of works develop neural networks in a Lagrangian framework [cranmer2019lagrangian, lutter2019deeplagrangiannetworksusing], which also preserve learned energies. More broadly, learning an energy and optimizing its gradient is an established framework in ML-based molecular dynamics simulation [GDML, gasteiger2022directionalmessagepassingmolecular, schütt2017schnetcontinuousfilterconvolutionalneural, zhang2018endtoendsymmetrypreservinginteratomic, Batzner_2022]. Despite research in this area, no prior works have considered learning with infinite-dimensional Hamiltonian systems and most HNNs have only been demonstrated in simple, particle-based systems.
Neural PDE surrogates are a growing field, with many works proposing architectures to improve model accuracy [lippe2023pderefinerachievingaccuratelong, wu2024transolverfasttransformersolver, alkin2025universalphysicstransformersframework], generalizability [zhou2024strategiespretrainingneuraloperators, hao2024dpotautoregressivedenoisingoperator, herde2024poseidonefficientfoundationmodels], or stability [zhou2025text2pdelatentdiffusionmodels, li2022learningdissipativedynamicschaotic, LIST2025117441, li2025generativelatentneuralpde]. Specific to our work, many PDE surrogates have also employed physics-based prior knowledge to improve model performance and adherence to governing laws. The most straightforward approach is to use physics-informed losses [RAISSI2019686, li2023physicsinformedneuraloperatorlearning, wang2021learningsolutionoperatorparametric], but another, more fundamental approach is to enforce physical symmetries and equivariances [shumaylov2025liealgebracanonicalizationequivariant, wang2021incorporating, brandstetter2023cliffordneurallayerspde, akhoundsadegh2023liepointsymmetryphysics, horie2023physicsembeddedneuralnetworksgraph]. Interestingly, by Noether’s theorem, every symmetry has an associated conservation law, which HNFs rely on as an inductive bias; through this lens the current work is another way of softly enforcing symmetries in predicted solutions, specifically time-translation symmetry. Despite previous works that incorporate physical priors into models, this work is likely the first to extend Hamiltonian mechanics to PDE surrogates, enabled by introducing neural functionals.
7 Conclusion
In this work, we have proposed a novel method for learning function to scalar mappings, enabling the extension of Hamiltonian mechanics to neural PDE surrogates. We demonstrate that neural functionals are able to implicitly learn functional derivatives, as well as achieve good performance on predicting 1D and 2D PDE solutions. In addition, Hamiltonian Neural Functionals have good capabilities in conserving energy-like quantities, enabling it to generalize to certain out-of-distribution inputs, such as a longer time horizon or unseen initial conditions. We hope that future work can continue to improve neural functionals as well as take advantage of functional approximation in new and insightful applications.
@@ 7 Conclusion
[…]
@@ Appendix A Additional Visualizations
Figure 5: Predicted trajectories of the KdV equation. The white dashed line denotes the training horizon, where only states from $t=0$ to $t=50$ are seen during training. Despite not seeing later, more chaotic states, HNFs still extrapolate beyond the training horizon by conserving the Hamiltonian, while other models struggle. FNO and Unet models are shown with their $d\mathbf{u}/dt$ variants.
@@
Figure 7: Predicted SWE trajectories with Gaussian pulse ICs, displayed at $t=10$ and $t=50$. Despite showing different behavior, HNFs can better generalize to OOD samples.
@@ B.1 Hamiltonian Errors and Inductive Bias
HNF 0.0002 1.32 0.0015 0.0003
Table 4: Relative L2 Error for each experiment, evaluated on the Hamiltonian of predicted trajectories. Despite not including the Hamiltonian in the training loss and testing on OOD samples, HNFs are exceptional at predicting solutions that conserve the Hamiltonian.
In the main body, we presented qualitative evidence that HNFs are able to better conserve the Hamiltonian of predicted trajectories. To quantify this claim, we calculate the Relative L2 Error of the Hamiltonian of predicted trajectories across all PDE experiments. Specifically, the error of a sample is given by: $\frac{1}{T}\sum_{t=1}^{T}\frac{||\mathcal{H}[\mathbf{u}^{t}]-\mathcal{H}[\mathbf{u}_{\theta}^{t}]||_{2}^{2}}{||\mathcal{H}[\mathbf{u}^{t}]||_{2}^{2}}$, where $\mathbf{u}^{t}$ is the true solution at time $t$, $\mathbf{u}^{t}_{\theta}$ is the predicted solution at time $t$, and $\mathcal{H}[\cdot]$ is the Hamiltonian given by the integrals in Equations 8 and 12. The Hamiltonian errors are averaged across the validation sets and reported in Table 4. We note that in the KdV experiments, the dynamics are chaotic and the Hamiltonian is highly nonlinear, both of which contribute to large errors once predicted trajectories diverge. Overall, we find that HNFs have exceptional capabilities in implicitly conserving energy-like quantities in their prediction.
@@
    ODE Bias: HNFs predict $\frac{d\mathbf{u}}{dt}$ and use an ODE integrator to evolve PDE dynamics.
@@
    Hamiltonian Bias: HNFs rely on $\mathcal{J}\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$ to calculate $\frac{d\mathbf{u}}{dt}$.
@@
    Gradient Learning Bias: HNFs rely on $\text{autograd}(\mathcal{H}_{\theta}[\mathbf{u}],\mathbf{u})$ to calculate $\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$.
@@
    Neural Functional Bias: HNFs rely on neural functionals to calculate $\text{autograd}(\mathcal{H}_{\theta}[\mathbf{u}],\mathbf{u})$.
@@
We find that the main contributors to HNF performance and its ability to conserve energies over time is using Hamiltonian structure as well as the neural functional architecture. Using an ODE integrator alone degrades model performance and implementing gradient learning with Unets also harms performance. Interestingly, using gradient learning with conventional architectures disproportionately harms energy conservation; this implies that in PDEs, gradient domain learning improves energy conservation only when architectures have meaningful functional derivatives. This is not readily apparent, as observations from molecular dynamics or HNNs suggest that gradient domain learning generally improves energy conservation without considering the underlying architectures [greydanus2019hamiltonianneuralnetworks, cranmer2019lagrangian, GDML]. Based on these findings, we hypothesize the following mechanism underlying HNFs: Neural functionals enable the learning of functional derivatives, combining this with infinite-dimensional Hamiltonian mechanics leads to its performance and conservation properties. Conveniently, these two contributions are the main novelties of this work.
@@
HNF 151.75 1.32
Table 5: Correlation time and Hamiltonian errors for Unet models with increasingly more inductive biases, compared to HNFs on the KdV equation. Using ODE integrators or gradient domain learning both degrade performance, while using Hamiltonian structure or neural functionals both increase performance and energy conservation.

### v2

@@ 4.1 Toy Examples
[…]
Table 2: Results for 1D PDEs. Parameter counts are: FNO (65K), Unet (65K), HNS (32K) for Adv, and FNO (135K), Unet (146K), HNS (87K) for KdV.
Figure 3: Hamiltonians for Adv and KdV for predicted trajectories, plotted over time. FNO and Unet models are shown with their $d\mathbf{u}/dt$ variants. HNS can better conserve Hamiltonians and remain stable over time.
Two baselines are considered, a FNO [13] and a Unet [25] model. Additionally, the HNS model is used with a nonlinear, FiLM-conditioned SIREN kernel. Global conditioning is used in the KdV experiments by calculating the scale/shift parameters as $\gamma_{i}^{(l)}=[\gamma_{\theta}^{(l)}(\mathbf{u}^{t})]_{i}$ and $\beta_{i}^{(l)}=[\beta_{\theta}^{(l)}(\mathbf{u}^{t})]_{i}$; the global FiLM networks are each parameterized by a shallow, 1D CNN. We find that global conditioning is needed in more complex PDEs since the kernel output $z_{i}=k_{\theta}(\cdot)$ can depend on more than just $u_{i}$ at a given point $x_{i}$. In particular, the Hamiltonian for the KdV equation has a $u_{xx}$ term that requires non-local knowledge of $\mathbf{u}^{t}$. Lastly, to ensure a fair comparison, we also train variants of the FNO and Unet baselines that predict temporal derivatives $\text{d}\mathbf{u}/\text{d}t$ [26], as opposed to the usual variants that predict $\mathbf{u}^{t+1}$; these use the same numerical integrator during inference as the HNS model.
[…]
We observe that the proposed HNS model is able to outperform current baselines, both on quantitative metrics and when conserving Hamiltonians over time. HNS is also efficient, capable of achieving better performance with nearly half the parameters of baseline models due to sharing weights between input points $x_{i},u_{i}$. Furthermore, despite training on a shorter time horizon, HNS can extrapolate and remain stable in longer prediction horizons. This is more pronounced in the KdV equation, where initial states are smoother than later, more chaotic states. An example trajectory is shown in Figure 5, where the baselines struggle to predict later states since these are not seen during training.
@@
To evaluate HNS on a more complex system, we consider the 2D Shallow Water Equations (SWE). These equations are an approximation of the Navier-Stokes equations when the horizontal length scale is much larger than the vertical length scale. We consider its conservative form, which is often used as a simple model for atmosphere or ocean dynamics.
For vector-valued functions $\mathbf{u}(x,y)\in\mathbb{R}^{d}$, the operator $\mathcal{J}$ becomes a $d\times d$ matrix of operators. For the 2D Shallow Water Equations, the operator matrix $\mathcal{J}$ is given by [27, 28, 29]:
@@
Data for the 2D SWE system is generated from PyClaw [30, 31, 32] using random sinusoidal initial conditions (Sines) for the height:
@@
[…]
Table 3: Rollout errors for 2D SWE. Parameter counts are: FNO/PINO (7M), Transolver (4M), Unet (3M), HNS (3M). Models are evaluated on in-distribution ICs (Sines) or out-of-distribution ICs (Pulse).
Figure 4: Hamiltonians for Sines and Pulse ICs, plotted over time. Despite potentially larger rollout errors, HNS can better conserve the Hamiltonian of predicted solutions.
@@
We consider benchmarking against Unet and FNO models, as well as PINO [33] and Transolver [34] models to compare against a physics-informed method and a newer attention-based surrogate. We instantiate the HNS model with a global, FiLM-conditioned SIREN kernel. Validation rollout errors for in-distribution ICs (Sines) and out-of-distribution ICs (Pulse) are given in Table 3, where means and standard deviations are calculated over three seeds. Additionally, the Hamiltonian over time for model rollouts are plotted in Figure 4. The derivative variants of baselines did not train stably, possibly due to sharp gradients that develop in the absence of viscosity, and are not reported. Additionally, the baselines are trained with the benefit of normalization, as the height channel has a different scale than the velocity channels; however, since the Hamiltonian is not invariant under scaling and shifting, normalization was not performed for the HNS model.
@@
For test samples within the training distribution (Sines), conventional neural surrogates are able to perform well and achieve low loss, and HNS is on-par with these models. Adding a physics-informed loss in PINO only marginally increases the performance of FNO, which is consistent with its interpretation as a soft regularizer. When looking at the Hamiltonian of predicted solutions, we find that HNS can predict solutions that better conserve physical invariances. We validate this in Appendix B.1 by calculating the error of the Hamiltonian of predicted trajectories across different experiments. This ability to conserve the Hamiltonian helps HNS to generalize to unseen initial conditions, and it achieves better performance on the Gaussian Pulse test set. Interestingly, it also conserves the Hamiltonian on out-of-distribution inputs, suggesting a strong inductive bias in the model. To understand what mechanisms enable energy conservation, we also ablate different methods within HNS to understand their contributions, shown in Appendix B.1. Lastly, to visualize model performance, predicted trajectories of Sines and Pulse ICs are shown in Figures 6 and 7.
5 Related Works
Hamiltonian Neural Networks
Original work on Hamiltonian Neural Networks (HNNs) [35, 36] established them as a promising architecture to respect physical laws, while later work improved their performance [37, 38, 39, 40, 41], investigated the mechanisms behind HNNs [42], or used HNNs in different applications [43, 44, 45]. A closely related set of works develop neural networks in a Lagrangian framework [46, 47], which also preserve learned energies. More broadly, learning an energy and optimizing its gradient is an established framework in ML-based molecular dynamics simulation [48, 49, 50, 51, 52]. Despite research in this area, no prior works have considered learning with infinite-dimensional Hamiltonian systems and most HNNs have only been demonstrated in simple, particle-based systems.
Functional Approximation
While uncommon, there are prior works that investigate learning infinite-dimensional to scalar/vector maps. Most methods are based on a kernel integral and are used to fit functionals [2], construct functional autoencoders [4], model parameter-to-observable maps [3], or build infinite-dimensional discriminators for generative modeling [5]. The use of a kernel integral to model a functional is also well studied in machine learning for DFT, since the Hohenberg-Kohn Theorems guarantee that the total energy is the integral of an unknown function of electron density [53, 54, 55]. Lastly, a separate approach considers modeling functionals through the cylindrical approximation, rather than a kernel integral [56].
Neural PDE Solvers
Neural PDE solvers are a growing field, with many works proposing architectures to improve model accuracy [57, 34, 58], generalizability [59, 60, 61], or stability [62, 63, 64, 65]. Specific to our work, many PDE solvers have also employed physics-based prior knowledge to improve model performance and adherence to governing laws. The most straightforward approach is to use physics-informed losses [66, 33, 67], but another, more fundamental approach is to enforce physical symmetries and equivariances in the network itself [68, 69, 70, 71, 72]. Interestingly, by Noether’s theorem, every symmetry has an associated conservation law, which HNS relies on as an inductive bias; through this lens the current work is another way of softly enforcing symmetries in predictions.
6 Conclusion
@@
Despite an interesting theoretical basis and its novelty, Hamiltonian Neural Solvers have limitations that we hope future work can address. Firstly, the theory for nonlinear functionals is underdeveloped. There is also additional overhead during inference to backpropagate through the network and evaluate the linear operator $\mathcal{J}$. Scalability is also a concern due to the reliance on numerical quadratures. This increase in runtime is quantified in Appendix B.2
Additionally, implementing $\mathcal{J}$ without a neural operator may require care to avoid numerical errors; this is further discussed in Appendix D.1. Furthermore, HNS is only applicable to Hamiltonian systems; this would exclude systems with dissipation, although modifications exist to extend Hamiltonian mechanics to non-conservative systems [73, 74, 75, 76]. Even for conservative PDEs, implementing a Hamiltonian structure is not as straightforward as training a neural solver with next-step prediction. Lastly, while not a limitation, an interesting observation is that introducing a loss on $\mathcal{H}[\mathbf{u}^{t}]-\mathcal{H}[\mathbf{u}_{\theta}^{t}]$ harms HNS performance; we hypothesize that since the Hamiltonian is conserved over time, learning the identity mapping is the easiest way to minimize this error and thus degrades model training.
@@
In this work, we have proposed a novel method for designing neural PDE solvers that respect conservation laws. We verify that kernel integrals are able to implicitly learn functional derivatives, as well as propose parametrization using neural fields. Using the resulting architecture in a Hamiltonian framework allows stable and energy-conserving predictions of 1D and 2D PDEs. These capabilities enable Hamiltonian Neural Solvers to generalize to certain out-of-distribution inputs, such as a longer time horizon or unseen initial conditions.
While learning functions and operators are dominant paradigms, approximating functionals may also have interesting uses. Functionals provide concise descriptions of physical systems by integrating physical variables, such as describing the energy of a molecule or the drag of an airfoil. Although concise, functional derivatives are often more relevant and can contain dynamical information or perhaps be used for optimization. We hope that future work can continue to improve functional approximators as well as take advantage of these models in new and insightful applications.
@@
We would like to thank Dr. Amit Acharya for his insightful discussions and help in conceptualizing this work.
@@ Appendix A Additional Visualizations
Figure 5: Predicted trajectories of the KdV equation. The white dashed line denotes the training horizon, where only states from $t=0$ to $t=50$ are seen during training. Despite not seeing later, more chaotic states, HNS still extrapolates beyond the training horizon by conserving the Hamiltonian, while other models struggle. FNO and Unet models are shown with their $d\mathbf{u}/dt$ variants.
@@
Figure 7: Predicted SWE trajectories with Gaussian pulse ICs, displayed at $t=10$ and $t=50$. Despite showing different behavior, HNS can better generalize to OOD samples.
@@ B.1 Hamiltonian Errors and Inductive Bias
HNS 0.0002 1.32 0.0015 0.0003
Table 4: Relative L2 Error for each experiment, evaluated on the Hamiltonian of predicted trajectories. Despite not including the Hamiltonian in the training loss and testing on OOD samples, HNS is exceptional at predicting solutions that conserve the Hamiltonian.
@@
Interestingly, this capability exists even when not trained with a loss on the Hamiltonian (i.e., $\mathcal{H}[\mathbf{u}^{t}]-\mathcal{H}[\mathbf{u}_{\theta}^{t}]||_{2}^{2}$) and when predicting out-of-distribution trajectories. Where does this inductive bias to conserve the Hamiltonian come from? Based on Gruver et al. [42], we hypothesize four inductive biases in HNS that may contribute to this:
@@
    ODE Bias: HNS predicts $\frac{d\mathbf{u}}{dt}$ and uses an ODE integrator to evolve PDE dynamics.
@@
    Hamiltonian Bias: HNS relies on $\mathcal{J}\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$ to calculate $\frac{d\mathbf{u}}{dt}$.
@@
    Gradient Learning Bias: HNS relies on $\text{autograd}(\mathcal{H}_{\theta}[\mathbf{u}],\mathbf{u})$ to calculate $\frac{\delta\mathcal{H}}{\delta\mathbf{u}}$.
@@
    Neural Functional Bias: HNS relies on integral kernel functionals to calculate $\text{autograd}(\mathcal{H}_{\theta}[\mathbf{u}],\mathbf{u})$.
@@
We find that the main contributors to HNS performance and its ability to conserve energies over time is using Hamiltonian structure as well as the neural functional architecture. Using an ODE integrator alone degrades model performance and implementing gradient learning with Unets also harms performance. Interestingly, using gradient learning with conventional architectures disproportionately harms energy conservation; this implies that in PDEs, gradient domain learning improves energy conservation only when architectures have meaningful functional derivatives. This is not readily apparent, as observations from molecular dynamics or HNNs suggest that gradient domain learning generally improves energy conservation without considering the underlying architectures [35, 46, 48]. Based on these findings, we hypothesize the following mechanism underlying HNS: integral kernel functionals enable the learning of functional derivatives, combining this with infinite-dimensional Hamiltonian mechanics leads to its performance and conservation properties.
@@
HNS 151.75 1.32
Table 5: Correlation time and Hamiltonian errors for Unet models with increasingly more inductive biases, compared to HNS on the KdV equation. Using ODE integrators or gradient domain learning both degrade performance, while using Hamiltonian structure or neural functionals both increase performance and energy conservation.